OpenStudy (anonymous):

What? a/-21 + 8 = 4 What's a? From my research, that answer would be 1.9047619, but I hear that's wrong....

OpenStudy (agreene):

\[\frac{a}{-21}+8=4\] is that what you're asking? or is it \[\frac{a}{-21+8}=4\]

OpenStudy (anonymous):

The former.

OpenStudy (anonymous):

And it's 2.1

OpenStudy (agreene):

\[\frac{a}{-2.1}+8=4\] \[-\frac{a}{2.1}=-4\] \[a=8.4\] \[a=\frac{42}{5}\]

OpenStudy (anonymous):

Thanks